OpenStudy (anonymous):

The rectangle shown has a length (l) of 8 and a width (w) of 6.2. A second rectangle has a length of l + 2.5 and a width of w – 1.4. Find the area of the SECOND rectangle.

OpenStudy (anonymous):

First find the dimensions you know l and w so the length of the second rectangle is length=(8)+2.5 and width=(6.2)-1.4